OpenStudy (anonymous):

Help please!? The sales tax on a new car varies directly to(or is directly proportional) the price of the car. If $1,800 in sales tax is paid for a $25,000 car. what is the purchase price of a car worth $1,600 in sales tax? Im totally lost when it comes to variations :(

OpenStudy (anonymous):

could you show the steps too?

OpenStudy (anonymous):

It's a question in proportions. 1800 is to 25000 as 1600 is to x. Are you with me so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but im not supposed to use proportions

OpenStudy (anonymous):

its for variations like this is a direct variation

OpenStudy (anonymous):

"varies directly" is the same as proportions. You will use an equation like a/b = c/x.

OpenStudy (anonymous):

1800 = a. 25000 = b. 1600 = c. Just get x

OpenStudy (anonymous):

25000/1800 = x/1600 25000/1800 = 13.89 1800 x 13.89 = 25 1600 x 13.89 = x

OpenStudy (anonymous):

the equation they gave to use to use was y=kx ^{n} \]

OpenStudy (anonymous):

y=kx squared to n

OpenStudy (anonymous):

What is your value for k?

OpenStudy (anonymous):

that what i dont know